OpenStudy (emilyjones284):

medal and fan! (: Evaluate the series 50 + 10 + 2 + . . .

jimthompson5910 (jim_thompson5910):

Hint: this is a geometric sequence with first term: a = 50 common ratio: r = 0.2

OpenStudy (primeralph):

@emilyjones284 I did something similar earlier.

jimthompson5910 (jim_thompson5910):

the common ratio was found by dividing any term by its previous term r = 10/50 = 1/5 = 0.2

OpenStudy (primeralph):